A gas has a volume of 17 liters at 20 oC. What will its volume at 100oC 


1
Expert's answer
2021-01-19T04:38:04-0500

V1=17 l, T1=20 oC, T2=100 oC. V2-?

V1/T1=V2/T2

17/293=V2/373

V2=21.64 l